# Active Learning and Explainable AI for Multi-Objective Optimization of Spin Coated Polymers

### v1

@@ 1 Introduction
Developing new materials is a challenging and time-consuming task. Traditionally, discovering new materials and optimizing synthesis procedures requires significant resources and long validation cycles (Collaert et al. 2019). One application where these challenges are noticeable is spin coating. Achieving an optimal spin-coated polymer film requires balancing multiple competing material properties. For example, both high hardness to resist deformation under load and low modulus of elasticity to remain flexible and to accommodate mechanical stress could be desired. Because materials design criteria are often contradictory, optimizing the resulting material becomes an extremely complex process, and global optimization is rarely possible. Instead, it is important to find design parameters that are non-dominated along particular criterion axes, the so-called Pareto front.
To address these challenges, AI is a powerful tool that can be leveraged to help accelerate materials discovery and optimization. These AI-driven methods can facilitate the identification of promising material compositions and guide the experimental process through active learning techniques to find the Pareto front. Bayesian optimization strategies (Erps et al. 2021), such as $\epsilon$-Pareto Active Learning ($\epsilon$-PAL) (Zuluaga, Krause, and Püschel 2016; Jablonka et al. 2021), have proven to be particularly effective in guiding experimental design by prioritizing the most informative combinations of parameters. However, when human researchers conduct these experiments, it is essential to provide transparent justifications for why certain parameter sets are selected and how it will benefit the materials science application.
@@
In this paper, we explore two different XAI techniques, fuzzy linguistic summaries (FLSs) (Kacprzyk and Zadrożny 2016) and Uniform Manifold Approximation and Projection (UMAP) (McInnes, Healy, and Melville 2020) methods, to generate explanations that support experimental justification in the context of optimizing spin-coated polyvinylpyrrolidone (PVP) films. By providing insights into the relationships between spin-coating parameters and material properties, these methods help researchers understand and validate AI-driven choices. Furthermore, when explanations are insufficient or domain expertise suggests alternative strategies, scientists can use these insights to refine and redirect experiments toward more promising regions of the design space. Figure 1 provides an overview of our process.
@@ 2.1 Related Work
This work intersects multiple concepts such as multi-criteria decision making (Emovon and Oghenenyerovwho 2020), optimization under uncertainty (Qian et al. 2023), design of experiments (Haridy, Gouda, and Wu 2011), and integration of domain knowledge with machine learning models to improve decision making and efficiency in materials discovery. Active learning has been applied to other areas of materials discovery, such as the optimization of chemical reaction conditions (Gong et al. 2021), the development of new battery materials (Verduzco, Marinero, and Strachan 2021), and the design of alloys (Xue et al. 2016). By combining these methodologies, AI can effectively navigate complex design spaces, identify optimal material compositions, and streamline the experimental process. Furthermore, these approaches enable the development of adaptive systems that learn from previous experiments and improve over time. It takes a combination of techniques such as these to overcome the numerous challenges in materials science, driving faster and more reliable advances in the creation of new materials with tailored properties.
@@ 2.2 $\epsilon$-PAL
An active learning technique designed for multi-objective optimization problems that we use is known as $\epsilon$-PAL (Zuluaga, Krause, and Püschel 2016). More specifically we are using a python implementation PyePAL (Jablonka et al. 2021). Traditional exhaustive search methods to identify the Pareto-optimal set of design parameters are computationally expensive, making active learning approaches like $\epsilon$-PAL highly beneficial. The core idea of $\epsilon$-PAL is to efficiently sample the design space while ensuring that the selected solutions approximate the true Pareto front within a user-defined tolerance, denoted by $\epsilon$. This approach allows users to sacrifice a specified amount of accuracy in order to significantly reduce the number of required experiments compared to conventional methods.
@@ 3.1 Fuzzy Linguistic Summaries
Although large language models (LLMs) have become incredible popular over the past few years with impressive products such as Open AI’s ChatGPT (OpenAI 2023), they cannot reliably analyze large datasets and often hallucinate answers (Nelson et al. 2024; Sriramanan et al. 2024). A more reliable family of techniques for producing accurate linguistic descriptions of data is known as fuzzy linguistic summaries (FLS). They can be useful for many things, including analyzing complex models that operate essentially as black boxes (Alvey, Anderson, and Keller 2023) (such as neural networks), for summarizing databases, (Niewiadomski 2008), explanations of time-series (Kacprzyk and Wilbik 2009), and anomaly detection (Wilbik and Keller 2013). Especially when considering several different variables at once, it can become nearly impossible to easily visualize the relationships between different combinations of conditions. This is where FLS excels, allowing for systematic evaluation of these combinations, determining which relationships hold true, and presenting them to the end-user.
@@
For each linguistic statement, various metrics can be computed. For example, truth (also known as validity) is a common metric which attempts to calculate how accurate a statement is with respect to a particular dataset. As noted in (Alvey, Anderson, and Keller 2025) and (Alvey et al. 2025), one must be careful to choose protoform structures which precisely corresponds to their respective metrics so that the values computed are meaningful. As such, we use the following equations for calculating truth values:
@@
These summaries can often result in many statements with high truth values that contain highly redundant information. For example, we might find that a polymer has near optimal material properties under a large number of different conditions. In this case, each condition, and combination of conditions would produce linguistic statements that evaluate to having high truth values. To address this redundancy we use a FLS simplification technique first presented in (Alvey, Anderson, and Keller 2025). This method involves representing the set of possible linguistic statements as a hierarchical directed acyclic graph which can be processed to determine which statements actually need to be reported. A user-defined threshold is employed to initially remove statements which do not have a high truth value. After the graph processing, the resulting intermediate FLS is passed to an LLM for further enhancement. This way, we take advantage of LLMs in a task they have been proven to be reliable in (summarizing and formatting), while leaving the bulk of information processing to a principled heuristic. This combination of modern AI and reliable pruning of the FLS results in preservation of important information presented in a much more user-friendly form, improving end-user insights and knowledge discovery.
@@ 3.2 UMAP
Uniform Manifold Approximation and Projection (UMAP) is a non-linear dimensionality reduction technique designed to preserve both global and local structures in high-dimensional data (McInnes, Healy, and Melville 2020). UMAP is grounded in manifold learning and graph theory, making it well-suited for visualizing complex datasets and uncovering underlying structures.

### v2

@@ 1 Introduction
Developing new materials is a challenging and time-consuming task. Traditionally, discovering new materials and optimizing synthesis procedures requires significant resources and long validation cycles (MaterialScienceHard). One application where these challenges are noticeable is spin coating. Achieving an optimal spin-coated polymer film requires balancing multiple competing material properties. For example, both high hardness to resist deformation under load and low modulus of elasticity to remain flexible and to accommodate mechanical stress could be desired. Because materials design criteria are often contradictory, optimizing the resulting material becomes an extremely complex process, and global optimization is rarely possible. Instead, it is important to find design parameters that are non-dominated along particular criterion axes, the so-called Pareto front.
To address these challenges, AI is a powerful tool that can be leveraged to help accelerate materials discovery and optimization. These AI-driven methods can facilitate the identification of promising material compositions and guide the experimental process through active learning techniques to find the Pareto front. Bayesian optimization strategies (3DPrinting), such as $\epsilon$-Pareto Active Learning ($\epsilon$-PAL) (ePAL; PyePAL), have proven to be particularly effective in guiding experimental design by prioritizing the most informative combinations of parameters. However, when human researchers conduct these experiments, it is essential to provide transparent justifications for why certain parameter sets are selected and how it will benefit the materials science application.
@@
In this paper, we explore two different XAI techniques, fuzzy linguistic summaries (FLSs) (FuzzySummaries1) and Uniform Manifold Approximation and Projection (UMAP) (UMAP) methods, to generate explanations that support experimental justification in the context of optimizing spin-coated polyvinylpyrrolidone (PVP) films. By providing insights into the relationships between spin-coating parameters and material properties, these methods help researchers understand and validate AI-driven choices. Furthermore, when explanations are insufficient or domain expertise suggests alternative strategies, scientists can use these insights to refine and redirect experiments toward more promising regions of the design space. Figure 1 provides an overview of our process.
@@ 2.1 Related Work
This work intersects multiple concepts such as multi-criteria decision making (MCDM), optimization under uncertainty (OUC), design of experiments (DoE), and integration of domain knowledge with machine learning models to improve decision making and efficiency in materials discovery. Active learning has been applied to other areas of materials discovery, such as the optimization of chemical reaction conditions (ChemicalReactions), the development of new battery materials (Batteries), and the design of alloys (Alloys). By combining these methodologies, AI can effectively navigate complex design spaces, identify optimal material compositions, and streamline the experimental process. Furthermore, these approaches enable the development of adaptive systems that learn from previous experiments and improve over time. It takes a combination of techniques such as these to overcome the numerous challenges in materials science, driving faster and more reliable advances in the creation of new materials with tailored properties.
@@ 2.2 $\epsilon$-PAL
An active learning technique designed for multi-objective optimization problems that we use is known as $\epsilon$-PAL (ePAL). More specifically we are using a python implementation PyePAL (PyePAL). Traditional exhaustive search methods to identify the Pareto-optimal set of design parameters are computationally expensive, making active learning approaches like $\epsilon$-PAL highly beneficial. The core idea of $\epsilon$-PAL is to efficiently sample the design space while ensuring that the selected solutions approximate the true Pareto front within a user-defined tolerance, denoted by $\epsilon$. This approach allows users to sacrifice a specified amount of accuracy in order to significantly reduce the number of required experiments compared to conventional methods.
@@ 3.1 Fuzzy Linguistic Summaries
Although large language models (LLMs) have become incredible popular over the past few years with impressive products such as Open AI’s ChatGPT (openai2023gpt4), they cannot reliably analyze large datasets and often hallucinate answers (nelson2024needlehaystackmemorybased; sriramanan2024llmcheck). A more reliable family of techniques for producing accurate linguistic descriptions of data is known as fuzzy linguistic summaries (FLS). They can be useful for many things, including analyzing complex models that operate essentially as black boxes (10195035) (such as neural networks), for summarizing databases, (4358814), explanations of time-series (5364881), and anomaly detection (6622481). Especially when considering several different variables at once, it can become nearly impossible to easily visualize the relationships between different combinations of conditions. This is where FLS excels, allowing for systematic evaluation of these combinations, determining which relationships hold true, and presenting them to the end-user.
@@
For each linguistic statement, various metrics can be computed. For example, truth (also known as validity) is a common metric which attempts to calculate how accurate a statement is with respect to a particular dataset. As noted in (Alvey2025_FUZZ_IEEE) and (Alvey2025_CAI), one must be careful to choose protoform structures which precisely corresponds to their respective metrics so that the values computed are meaningful. As such, we use the following equations for calculating truth values:
@@
These summaries can often result in many statements with high truth values that contain highly redundant information. For example, we might find that a polymer has near optimal material properties under a large number of different conditions. In this case, each condition, and combination of conditions would produce linguistic statements that evaluate to having high truth values. To address this redundancy we use a FLS simplification technique first presented in (Alvey2025_FUZZ_IEEE). This method involves representing the set of possible linguistic statements as a hierarchical directed acyclic graph which can be processed to determine which statements actually need to be reported. A user-defined threshold is employed to initially remove statements which do not have a high truth value. After the graph processing, the resulting intermediate FLS is passed to an LLM for further enhancement. This way, we take advantage of LLMs in a task they have been proven to be reliable in (summarizing and formatting), while leaving the bulk of information processing to a principled heuristic. This combination of modern AI and reliable pruning of the FLS results in preservation of important information presented in a much more user-friendly form, improving end-user insights and knowledge discovery.
@@ 3.2 UMAP
Uniform Manifold Approximation and Projection (UMAP) is a non-linear dimensionality reduction technique designed to preserve both global and local structures in high-dimensional data (UMAP). UMAP is grounded in manifold learning and graph theory, making it well-suited for visualizing complex datasets and uncovering underlying structures.